If the odds in favor of Fast Leg winning a horse race are 7 to 3 and the first prize is $20,000, what is the expected value of Fast​ Leg's winnings?
Answer #1

Ans:

P(Fast Leg's winnings)=7/(3+7)=7/10=0.7

Expected value of Fast​ Leg's winnings=0.7*20000=14000